OpenStudy (swagster23780):

What is the product in lowest terms? 5/6 x -8/15 A. B. C. D.

OpenStudy (swagster23780):

@mathmale @hoots89 PLEASE HELP

OpenStudy (swagster23780):

@Loser66 Help

OpenStudy (mathmale):

For clarity please use parentheses: 5/6 x -8/15 ambiguous (5/6)x - (8/15) clear

OpenStudy (mathmale):

We are to find the PRODUCT of (5/6)x and (-8/15). You'd probably find this to be much easier if you'd please write the product as follows: \[\frac{ 5 }{ 6 }*(\frac{ 8 }{ 15 })\]

OpenStudy (mathmale):

Note: Best if you use " * " to indicate multiplication. I wqs moving too fast and mistook your " x " for the letter x, when you meant multiplication.

OpenStudy (mathmale):

How would you multiply \[\frac{ 5 }{ 6 }*\frac{ 8 }{ 15 }\]

OpenStudy (mathmale):

???

OpenStudy (mathmale):

Hint: can 8/6 be reduced? Can 5/15 be reduced?

OpenStudy (swagster23780):

Idk Know Its Confusing

OpenStudy (mathmale):

please reduce this:\[\frac{ 8 }{ 6 }\]

OpenStudy (mathmale):

Also please reduce this: \[\frac{ 5 }{ 15 }\]

OpenStudy (swagster23780):

Please Just Tell Me

OpenStudy (swagster23780):

What It Is

OpenStudy (swagster23780):

@mathmale